Status Report

NASA Space Station On-Orbit Status 3 Jul 2004

By SpaceRef Editor
July 3, 2004
Filed under , , ,
NASA Space Station On-Orbit Status 3 Jul 2004

SpaceRef note: This NASA Headquarters internal status report, as presented here, contains additional, original material produced by (copyright © 2004) to enhance access to related status reports and NASA activities.

All ISS systems continue to function nominally except those noted previously or below.  The crew is enjoying a three-day weekend in celebration of our Independence Day holiday.

Sleep/wake cycle has return to normal, with wakeup this morning at 2:00am EDT.

After breakfast, Padalka and Fincke performed the regular weekly 3-hr. station cleaning.  [“Uborka” includes removal of food waste products, cleaning of compartments with vacuum cleaner, wet cleaning of the Service Module (SM) dining table and other surfaces with “Fungistat” disinfectant and cleaning fan screens to avoid temperature rises.]

At 9:30am, the crew conducted the weekly planning conference (WPC) with the ground, discussing next week’s “Look-Ahead Plan” (regularly prepared jointly by MCC-H and MCC-M planners), via S-band/audio.

For his “Saturday Science” program today, Science Officer Fincke performed another BCAT-3 (Binary Colloid Alloy Test 3) photography session.   [After setting up the SGSM (slow growth sample module) B at the Lab MWA (Maintenance Work Area), Mike conducted the videorecorded demo by photographing samples 4, 5, 6, 8, 9, and 10 with the Kodak 760 digital still camera, using an uplinked new procedure for fine focusing.  The images were then stored on a PCMCIA 1GB micro drive.  Mike stowed the samples and afterwards disassembled and also stowed the BCAT MWA hardware.]

Previous Reports

ISS On-orbit Status [HQ]

ISS Status [JSC]

Shuttle Processing [KSC]

The CDR set up the Orlan battery-charging unit (ZU-S), then initiated discharge of the two 825M3 Orlan backpack batteries for storage.  To allow monitoring of the procedure from the ground, Padalka first hooked up the Service Module (SM)’s BITS2-12 onboard telemetry system.

Afterwards, on “Go” from TsUP, Gennady downloaded the stored data of last night’s tests of the Russian ASN-M satellite navigation system’s NPM receiver module from Laptop 3 to a FlashCard for transmittal to the ground via the SAM computer system.

Later, working off the Russian discretionary task list, Padalka downloaded new accumulated Matryoshka measurement tables from the Matryoshka server (BSPN) via the ISS Wiener laptop to a PCMCIA memory card for subsequent downlink on OCA comm.   [Matryoshka automatically takes measurements in the SM and DC-1 docking compartment for studies of on-orbit radiation and long-term dose accumulation, using six SPD dosimeters deployed throughout the Russian segment as well as in a spherical body-simulating Matryoshka-R “phantom” and a human torso model outside on the SM hull, mounted there during EVA-9.]

Also on Padalka’s Russian task list for today was the downloading of the data from the latest run of the Molniya-SM/LSO experiment (6/20-23), from the LSO memory to HDD (hard-disk drive) on the EGE-1 laptop for return to Earth.   [Objective of the unattended Molniya-SM sessions, similar to the French LSO experiment, is to record storm phenomena and other related events in the Earth’s equatorial regions.]

ISS Location NOW

Full Size/Update
Real Time ISS TrackerMore Links

Gennady completed the regular daily maintenance/inspection of the SM’s SOZh environment control & life support systems, including the ASU toilet system.

New pages with SODF (operations data files) updates on emergency/warning procedures were uplinked yesterday, for Mike to print out and incorporate in the three Warning Books on board.

Yesterday at 1:33pm, the Russian SM TsVM (Central Computer) lost the second of its three redundant subsets (“lanes”) when string #1 was voted out of the set (string #3 had dropped out on 10/22/03).   [Impact: Reduced redundancy on SM TsVM.  A TsVM restart, which is required to integrate the two lanes back into the redundant set, is already scheduled for 7/14 to prepare for the SM 7.02 flight software upgrade.  There are currently no plans for an earlier restart.  If string #2 goes also down before that date, there would be a temporary loss (~2 min) of communication between RS and USOS systems until the processors are reintegrated by automatic reboot.]

[Previous NASA ISS On-orbit Status Reports can be found here. Previous NASA Space Station Status Reports can be found here. Previous NASA Space Shuttle Processing Status Reports can be found here. A collection of all of these reports and other materials relating to Return to Flight for the Space Shuttle fleet can be found here.]

In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

Weekly Science Update (Expedition Nine — 10th):

GASMAP:   The GASMAP Functional Check this week went flawlessly.  The next activity will be a Routine Health Check sometime next month.

Human Research Facility/Workstation (HRF WS):   Continuing.

Advanced Ultrasound (ADUM):   The ADUM team is eagerly looking forward to next week’s bone and abdominal scans.

Hand Posture Analyzer (HPA:  Nothing new.

In-Space Soldering Investigation (ISSE):   Nothing new.

Pore Formation and Mobility Investigation (PFMI):   Nothing new.
Space Acceleration Measurement System (SAMS):   A software modification is being prepared that will allow the SAMS ICU laptop to operate with a missing, defective, or discharged battery. The ground team will use an Express FTP process to load the new software. Initial software testing successfully completed at MSFC.
Microgravity Acceleration Measurement System (MAMS):   The MAMS OSS (Orbital Acceleration Research Experiment Sensor Subsystem) continues collection of quasi-steady acceleration measurements from the station.  The HiRAP (High Resolution Accelerometer Package) is also enabled to capture vibratory data below 100 Hz in EXPRESS Rack 1

Protein Crystal Growth-Single Locker Thermal Enclosure System (PCG-STES):  Nominal.

Protein Crystal Growth Monitoring by Digital Holographic Microscope (PromISS):   Nothing new.

Investigating the Structure of Paramagnetic Aggregates from Colloidal Emulsions (InSPACE):    Planned.

Binary Colloidal Alloy Test-3 (BCAT-3):   Complete.

Renal Stone (RS):   Nothing new.

Synchronized Position Hold, Engage, Reorient, Experimental Satellites (SHERES):   Nothing new.

Foot/Ground Reaction Forces during Space Flight (FOOT):   Nothing new.

Materials ISS Experiment (MISSE):  In progress. Deployed outside on the U.S. Airlock. Nominal and collecting data.

Cellular Biotechnology Support Systems-Fluid Dynamics Investigation (CBOSS-FDI):   Nothing new.

Biotechnology Specimen Temperature Controller (BSTC):   Planned.

Yeast Group Activation Packs (Yeast GAP):   Nothing new.

Earth Knowledge Acquired by Middle School Students (EarthKAM):  There is one more session scheduled for this Increment, in July.

Earth Science Toward Exploration Research (ESTER):  Nothing new.

Serial Network Flow Monitor (SNFM):   Planning run #3 on 7/8 using Configuration 2A (LAN-1).

Viscous Liquid Foam–Bulk Metallic Glass (Foam):  The Foam team members appreciate the crew’s choosing a conference with them as “Saturday Science”, and are looking forward to it and to operations.  This very foam was voted “Best of What’s Next 2004” by Popular Science (the team affectionately calls it “bubbloy”, derived from the words bubble and alloy).

BIOPSY (Effects of Prolonged Space Flight on Human Skeletal Muscle):  Nothing new.

Coarsening in Solid-Liquid Mixtures-2 (CSLM-2):  Planned.

Commercial Generic Bioprocessing Apparatus (CGBA):   Nothing new.

Miscible Fluids in Microgravity (MFMG):    Nothing new.

Educational Payload Operations (EPO):   In planning.

Crew Earth Observations (CEO):   Review of ISS/CEO imagery has been hampered the past ten days by a hard drive crash on the team’s main server.  However, it is hoped that responses to the crew’s most helpful CEO logs can be resumed early next week.  A colorful image of the Colorado River delta was published on NASA’s Earth Observatory website this week.  It highlights an article describing the delta’s delicate and changing environmental conditions due to a dwindling supply of fresh water in the river reaching the sea in a harsh, desert climate.

Today’s optional CEO photo targets, in the current LVLH attitude no longer limited by flight rule constraints on the use of the Lab nadir/science window, except for the shutter closure and condensation-prevention plan (limited to 90 min. in 24 hours), were Saharan dust, Morocco (Dynamic event.  Dust is streaming westward off the coast of Morocco, well north of the usual outflow position near the latitude of Gibraltar. Looking half left for views of the dust margin with the coast of Africa in the background), Albuquerque, New Mexico (nadir pass), and Iguazu, Brazil (the Iguazu Falls are the second largest tourist attraction in South America.  In the last decade the city nearby has become the fastest growing urban region in South America since it lies at the meeting point between Brazil, Argentina and Paraguay where trade is vigorous.  Looking a touch left at the confluence of the big Parana River and the tributary Iguazu River.  A prominent protected forest zone and major lake on the Parana are other pointers).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

To view the latest photos taken by the expedition 9 crew visit

U.S. & Russian Segment Status  (as of today, 1:45pm EDT)

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is On.  Vozdukh CO2 scrubber is On.  U.S. CDRA CO2 scrubber is Off.  TCCS (trace contaminant control subsystem) is operating.  SM Gas Analyzer has been calibrated and is used for ppO2 & ppCO2 monitoring.  MCA (major constituents analyzer) is in Life Extending Mode (LEM).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is Off, SKV-2 is Off (SM panel mods completed 4/8; SKV-2 activation failed 4/20; is now functioning again).  SFOG slot #2 fan suspect (not usable).

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in directed position (dual angle/”blind” mode, non solar-tracking, biased for drag reduction).
  • SM batteries:  All batteries (8) are on line in “Partial Charge” mode.  
  • FGB batteries:  Battery #5 is off line; all other batteries (5) are on line in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 is in Standby mode; PCU-2 is in Standby mode.

Command & Data Handling Systems:

  • C&C-1 MDM is prime, C&C-2 is backup, and C&C-3 is in standby.
  • GNC-2 MDM is prime; GNC-1 is backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-1 is On (primary), EXT-2 is Off (backup).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational (string 1 dropped out 11/22/03).
  • SM Central Computer (TsVM): 1 redundant lane (of 3) operational (string #3 dropped out 10/22/03; string #1 dropped out 7/2).
  • FGB MDM-1 is Off (failed, 11/21/03); MDM-2 is Operational.

Propulsion System:

  • Total propellant load available: 3928 kg (8660 lb) as of 7/1/04;  [SM(552) + FGB(2772) + Progress M(639)].  (Capabilities: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed, since 6/6/02; CMG-2’s RPC-17 failed 4/21/04; was replaced 6/30/04).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH XVV (local vertical/local horizontal = “earth-fixed”: z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -9 deg, roll: 0 deg]), with CMG/TA (thruster assist) Momentum Management, until 6/28, following the EVA.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally (may require a mask).
  • Audio subsystem is operating nominally (IAC-1 is prime, IAC-2 is off).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.


  • SSRMS/Canadarm2 based at Lab PDGF/LEE A, operational on redundant string, off on prime.
  • MBS: KA (keep alive) power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of this morning, 8:40am EDT [= epoch]):

  • Mean altitude — 360.1 km
  • Apogee — 363.8 km
  • Perigee — 356.4 km
  • Period — 91.7 min.
  • Inclination (to Equator) —  51.6321 deg
  • Eccentricity — 0.0005484
  • Solar Beta Angle — 44.0 deg
  • Orbits per 24-hr. day — 15.69
  • Mean altitude loss in last 24 hours — 65 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 32088

ISS Altitude History

Apogee height Mean AltitudePerigee height

ISS Altitude History

For more on ISS orbit and worldwide ISS naked-eye visibility dates/times, see In addition, information on International Space Station sighting opportunities can be found at on NASA’s Human Spaceflight website. The current location of the International Space Station can be found at at NASA’s Marshall Space Flight Center. Additional satellite tracking resources can be found at

SpaceRef staff editor.